# Michel F. C. Haddad

### v1

@@ 3 Identification in panel data
This section discusses identification in panel data under the identifying assumptions proposed in the previous section. We first reconsider the identification of $\Delta_{d_{t},d_{t}^{\prime},t}$, i.e., the ATET of treatment dose $d_{t}$ versus $d_{t}^{\prime}$ in period $t$ (net of direct effects of previous treatment doses) as defined in equation (7). As panel data permits taking outcome differences within subjects over time, the identification result in equation (2) simplifies as follows:
$\Delta_{d_{t},d_{t}^{\prime},t}$ $=$ $E[E[Y_{t}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]-E[Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$ (13)
-- $\{E[Y_{t}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$
-- $E[Y_{t-1}|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]\}|D_{t}=d_{t},T=t]$
$=$ $E[E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$
@@
It is worth noticing that dropping $T=t$ from the conditioning set after the second equals sign follows from the fact that in panel data, the same subjects are observed in periods $t$ and $t-1$, such that distributions of $\mathbf{D_{t-1}}$ and $X$ are constant across periods by design.
To ease notation, we henceforth denote conditional mean differences within treatment groups across time periods by $m_{d_{t}}(t,\mathbf{D_{t-1}},X)=E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}},X]$. Furthermore, we denote by $P_{d}$ and $p_{d,t}(\mathbf{D_{t-1}},X)$ (conditional) density functions of treatment $D=d$ that correspond to $P_{d_{t}}=\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)]$ and $p_{d_{t}}(\mathbf{D_{t-1}},X)=\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)|\mathbf{D_{t-1}},X]$, respectively. We propose an DR expression that is a straightforward modification of the one provided in equation (3.8) of Zhang (2024). The key differences are that the dose in the control group is non-zero - thus, involving kernel functions with respect to both $d_{t}^{\prime}$ and $d_{t}$, and that the conditioning set includes the treatment history $\mathbf{D_{t-1}}$ due to the multiple-period setup. Our proposed DR expression is equivalent to equation (13) for ATET identification, being given by:
@@
$-\left.\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D_{t-1}},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D_{t-1}},X)]}{p_{d_{t}^{\prime}}(\mathbf{D_{t-1}},X)\cdot P_{d_{t}}}\right]$ (14)
To see the equivalence, we note that by the law of iterated expectation,
@@
which corresponds to the ATET of interest. In addition, applying basic probability theory and the law of iterated expectations, we have that
@@
which demonstrates that both equations (3) and (13) identify the ATET under Assumptions 4 to 7.
Lastly, we consider evaluating $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$, the ATET of receiving the treatment dose $d$ versus $d^{\prime}$ in period $t-s$ on a later outcome in period $t$, as defined in equation (11). Under the satisfaction of Assumptions 4, 6, 7, and 8, the DR expression (3) is modified to the following equation:
@@
where $m_{d_{t-s}^{\prime}}(t,\mathbf{D_{t-s-1}},X)=E[Y_{t}-Y_{t-s-1}|D_{t-s}=d_{t-s},\mathbf{D_{t-s-1}},X]$. Appendix A.1 demonstrates that DR expressions (3), and (15) satisfy Neyman orthogonality.
@@ 4 Estimation
In this section, we propose DiD estimators based on DML framework with cross-fitting, as discussed in Chernozhukov et al. (2018), which are based on the sample analogs of the DR identification results in sections 2 and 3. We first consider estimating the ATET $\Delta_{d_{t},d_{t}^{\prime},t}$ as defined in equation (7) based on the sample analog of equation (2). Thus, let $\mathcal{W}=\{W_{i}|1\leq i\leq n\}$ with $W_{i}=(Y_{i,T},D_{i,t},\mathbf{D_{i,t-1}}X_{i},T_{i})$ for $i=1,\ldots,n$ denote the set of observations in an i.i.d. sample of size $n$. Moreover, let $\mathcal{W}=\{W_{i}|1\leq i\leq n\}$ with $W_{i}=(Y_{i,T},D_{i,t},\mathbf{D_{i,t-1}}X_{i},T_{i})$ for $i=1,\ldots,n$ denote the set of observations in an i.i.d. sample of size $n$. Furthermore,
@@
denotes the set of nuisance parameters, i.e., the conditional mean outcomes and treatment densities. Their respective estimates are denoted by
@@
In order to avoid overfitting, we suggest estimating $\Delta_{d_{t},d_{t}^{\prime},t}$ based on the following DML algorithm with cross-fitting, where the (kernel- and generalized propensity score-related) weights within each group are defined upon the treatment state, and the time period are normalized to sum up to one:
@@
    $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}=\sum_{i=1}^{n}\left[\biggl{(}\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}\biggr{)}^{-1}\cdot(\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}\right.$
@@
    $-\frac{\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})\cdot\hat{\epsilon}_{i,d_{t},t-1,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d_{t},t-1}(\mathbf{D_{i,t-1}},X)}{\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d_{t},t-1}(\mathbf{D_{i,t-1}},X)}$
    $-\left(\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D_{i,t-1}},X_{i})}\right.$
    $-\left.\left.\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t-1,\mathbf{D_{i,t-1}},X_{i}}/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D_{i,t-1}},X_{i})}\right)\right],$
@@
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ $=\sum_{i=1}^{n}\left[\frac{\omega(D_{i,t};d_{t},h)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D_{i,t-1}},X_{i})]}{\sum_{i=1}^{n}\omega(D_{i,t};d_{t},h)}\right.$
$\left.-\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot\hat{\rho}_{d_{t}}(\mathbf{D_{i,t-1}},X)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D_{i,t-1}},X_{i})]/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D_{i,t-1}},X_{i})}{\sum_{i=1}^{n}\omega(D_{i,t};d^{\prime}_{t},h)\cdot\hat{\rho}_{d_{t}}(\mathbf{D_{i,t-1}},X_{i})/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D_{i,t-1}},X_{i})}\right].$
Furthermore, the algorithm may be easily adapted to estimating $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$, the ATET of a specific treatment dose of an earlier period on a later outcome as defined in (11), by using of the normalized sample analogs of DR expressions (2) or (15) in the case of repeated cross-sections or panel data, respectively.
For estimating the variance of our estimators to conduct inference, we adapt the asymptotic variance approximation suggested in Zhang (2024) to our scenarios with non-zero treatment doses, in both the treatment and control groups. In the case of repeated cross-sections, the asymptotic variance of the ATET estimator of $\Delta_{d_{t},d_{t}^{\prime},t}$ given in (7) based on the sample analog of equation (2), denoted by $\sigma_{h}^{2}$, corresponds to
Defining
@@
we have that $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E\left[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)\right]$ in the variance formula, which is the ATET approximation using bandwidth $h$, while $\epsilon_{{d_{t}},t,\mathbf{d_{t-1}},x}=(Y_{T}-\mu_{d_{t}}(t,\mathbf{d_{t-1}},x))$ is a regression residual. Furthermore, $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the Neyman-orthogonal score function under bandwidth $h$. It is worth noticing that the second term in the asymptotic variance in equation (16)
@@
comes for a Taylor expansion of the score w.r.t. the (low dimensional) density $P(D_{t}=d_{t},T=t)$. Therefore, it reflects estimation uncertainty related to the kernel approach required for continuous treatment doses.
A variance estimator may be constructed based on the sample analog of equation (16):
$\hat{\sigma}_{h}^{2}=\frac{1}{n}\sum_{i=1}^{n}\left[\left(\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta})-\frac{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}}{\hat{\Pi}_{k,d_{t},t}}\left(\omega(D;d,h)\cdot I\{T=t\}-\hat{\Pi}_{k,d_{t},t}\right)\right)^{2}\right],$ (20)
where $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ is the cross-fitted ATET estimate (see step 4 of the algorithm). The term $\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta})$ denotes the estimate of score function $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)$ for subject $i$, which is computed using the cross-fitted nuisance parameters $\hat{\eta}$ (see step 3 of the algorithm). $\hat{\Pi}_{d_{t},t}=\frac{1}{n}\sum_{i=1}^{n}\omega(D_{i};d,h)\cdot I\{T_{i}=t\}$ an estimate of the density $\Pi_{d_{t},t}$. In Zhang (2024), such types of variance estimators are shown to be consistent, provided that the kernel bandwidth $h$ satisfies $h^{-2}\varepsilon_{n}^{2}=o(1)$ for $\varepsilon_{n}=o(n^{-1/4})$.
@@
where $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]$ is the ATET approximation under bandwidth $h$ and $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the score function, with

### v2

@@ 4 Identification in panel data
This section discusses identification in panel data under the identifying assumptions proposed in the previous section. We first consider the identification of $\Delta_{d_{t},d_{t}^{\prime},t}$, i.e., the ATET of treatment dose $d_{t}$ versus $d_{t}^{\prime}$ in period $t$ (net of direct effects of previous treatment doses) as defined in equation (6). As panel data permits taking outcome differences within subjects over time, the identification result in equation (3) simplifies as follows:
$\Delta_{d_{t},d_{t}^{\prime},t}$ $=$ $E[E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D}_{t-1}=\mathbf{D}_{t-1},X=x]$ (18)
@@
It is worth noticing that dropping $T=t$ from the conditioning follows from the fact that, in panel data, the same subjects are observed in periods $t$ and t1t-1, such that distributions of $\mathbf{D}_{t-1}$ and $X$ are constant across periods by design. To ease notation, we denote the conditional mean differences in outcomes by $m_{d_{t}}(t,\mathbf{D}_{t-1},X)=E[Y_{t}-Y_{t-1}|D_{t}=d_{t},\mathbf{D}_{t-1},X]$. Moreover, we use $P_{d_{t}}=f(D_{t}=d_{t})$ and $p_{d_{t}}(\mathbf{D}_{t-1},X)=f(D_{t}=d_{t}|\mathbf{D}_{t-1},X)$ to denote the (conditional) density functions of treatment $D_{t}=d_{t}$. Specifically, it can be shown that $P_{d_{t}}=\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)]$ and $p_{d_{t}}(\mathbf{D}_{t-1},X)$$\lim_{h\rightarrow 0}E[\omega(D;d_{t},h)|\mathbf{D}_{t-1},X]$ hold. We propose an DR expression that is a straightforward modification of the one provided in equation (3.8) of Zhang (2024). The key differences are that the treatment dose in the control group is non-zero, thus involving kernel functions with respect to both $d_{t}^{\prime}$ and $d_{t}$, and that the conditioning set additionally includes the treatment history $\mathbf{D}_{t-1}$ due to the multiple-period setup. Our proposed DR expression for equation (18) for ATET identification is given by:
@@
$-\left.\frac{\omega(D_{t};d_{t}^{\prime},h)\cdot p_{d_{t}}(\mathbf{D}_{t-1},X)\cdot[Y_{t}-Y_{t-1}-m_{d_{t}^{\prime}}(t,\mathbf{D}_{t-1},X)]}{p_{d_{t}^{\prime}}(\mathbf{D}_{t-1},X)\cdot P_{d_{t}}}\right].$ (19)
To see the equivalence of equations (4) and (18), we note that by the law of iterated expectation,
@@
which corresponds to the ATET of interest. In addition, again by the law of iterated expectations, we have that
@@
which demonstrates that both equations (4) and (18) identify the ATET under Assumptions 4 through 7 (without conditioning on $T=t$).
Next, we consider evaluating $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$, the ATET of receiving the treatment dose $d$ versus $d^{\prime}$ in period tst-s on a later outcome in period $t$, as defined in equation (10). Under Assumptions 4, 6, 7, and 8, a DR expression in the panel setting with time lags, analogous to (3) for repeated cross sections, is given by the following expression:
@@
where $m_{d_{t-s}^{\prime}}(t,\mathbf{D}_{t-s-1},X)=E[Y_{t}-Y_{t-s-1}|D_{t-s}=d_{t-s},\mathbf{D}_{t-s-1},X]$. Appendix A.1 demonstrates that DR expressions (4) and (20) satisfy Neyman orthogonality.
Finally, when assessing the ATET $\Delta_{d_{t-s},\mathbf{d^{\prime}}_{t-s:t},t}$ under the fixed treatment sequence defined in equation (14), equation (20) needs to be modified by making use of the product kernel defined in equation (16), the conditional mean outcome differences under fixed treatment sequences, $m_{\mathbf{d}_{t-s:t}^{\prime}}(t,\mathbf{D}_{t-1},X)=E[Y_{t}-Y_{t-1}\mid\mathbf{D}_{t-s:t}^{\prime}=\mathbf{d}_{t-s:t}^{\prime},\mathbf{D}_{t-1},X]$, and the corresponding conditional densities of fixed treatment sequences, $p_{\mathbf{d}_{t-s:t}^{\prime}}(\mathbf{D}_{t-1},X)=f(\mathbf{D}_{t-s:t}^{\prime}=\mathbf{d}_{t-s:t}^{\prime}\mid\mathbf{D}_{t-1},X)$. Under Assumptions 4, 6, 9, and 10, the following expression identifies the ATET in panel data:
$\Delta_{d_{t-s},\mathbf{d}_{t-s:t}^{\prime},t}$ $=\lim_{h\rightarrow 0}E\left[\frac{\omega(D_{t-s};d_{t-s},h)\cdot[Y_{t}-Y_{t-s-1}-m_{\mathbf{d}_{t-s:t}^{\prime}}(t,\mathbf{D}_{t-s-1},X)]}{P_{d_{t-s}}}\right$. (21)
$-\left.\frac{\omega(\mathbf{D}_{t-s:t};\mathbf{d}^{\prime}_{t-s:t},\mathbf{h})\cdot p_{d_{t-s}}(\mathbf{D}_{t-s-1},X)\cdot[Y_{t}-Y_{t-s-1}-m_{\mathbf{d}_{t-s:t}^{\prime}}(t,\mathbf{D}_{t-s-1},X)]}{p_{\mathbf{d}_{t-s:t}^{\prime}}(\mathbf{D}_{t-s-1},X)\cdot P_{d_{t-s}}}\right].$
@@ 5 Estimation
In this section, we propose DiD estimators based on the DML framework as discussed in Chernozhukov et al. (2018), using the sample analogs of the DR identification results in sections 2 and 4. We first consider estimating the ATET $\Delta_{d_{t},d_{t}^{\prime},t}$ as defined in equation (6) with the sample version of equation (3). Thus, let $\mathcal{W}=\{W_{i}|1\leq i\leq n\}$ with $W_{i}=(Y_{i,T},D_{i,t},\mathbf{D}_{i,t-1}X_{i},T_{i})$ for $i=1,\ldots,n$ denote the set of observations in an i.i.d. sample of size $n$. Moreover, let
@@
denote the set of nuisance parameters, i.e., the conditional mean outcomes and treatment densities. Their respective estimates are denoted by
@@
The following algorithm formally sumarizes the estimation procedure for $\Delta_{d_{t},d_{t}^{\prime},t}$:
@@
    $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ $=\sum_{i=1}^{n}\left[\biggl(\sum_{j=1}^{n}\omega(D_{j,t};d_{t},h)\cdot I\{T_{j}=t\}\biggr)^{-1}\cdot(\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t\}\right$.
@@
    $-\frac{\omega(D_{i,t};d_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X_{i})\cdot\hat{\epsilon}_{i,d_{t},t-1,\mathbf{D}_{i,t-1},X_{i}}/\hat{\rho}_{d_{t},t-1}(\mathbf{D}_{i,t-1},X)}{(\sum_{j=1}^{n}\omega(D_{j,t};d_{t},h)\cdot I\{T_{j}=t-1\})\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X_{i})/\hat{\rho}_{d_{t},t-1}(\mathbf{D}_{i,t-1},X)}$
    $-\left(\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t,\mathbf{D}_{i,t-1},X_{i}}/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D}_{i,t-1},X_{i})}{(\sum_{j=1}^{n}\omega(D_{j,t};d^{\prime}_{t},h)\cdot I\{T_{j}=t\})\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X_{i})/\hat{\rho}_{d^{\prime}_{t},t}(\mathbf{D}_{i,t-1},X_{i})}\right$.
    $-\left.\left.\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot I\{T_{i}=t-1\}\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X)\cdot\hat{\epsilon}_{i,d^{\prime}_{t},t-1,\mathbf{D}_{i,t-1},X_{i}}/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D}_{i,t-1},X_{i})}{(\sum_{j=1}^{n}\omega(D_{j,t};d^{\prime}_{t},h)\cdot I\{T_{j}=t-1\})\cdot\hat{\rho}_{d_{t},t}(\mathbf{D}_{i,t-1},X_{i})/\hat{\rho}_{d_{t}^{\prime},t-1}(\mathbf{D}_{i,t-1},X_{i})}\right)\right],$
@@
$\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ $=\sum_{i=1}^{n}\left[\frac{\omega(D_{i,t};d_{t},h)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D}_{i,t-1},X_{i})]}{\sum_{j=1}^{n}\omega(D_{j,t};d_{t},h)}\right$.
$\left.-\frac{\omega(D_{i,t};d^{\prime}_{t},h)\cdot\hat{\rho}_{d_{t}}(\mathbf{D}_{i,t-1},X)\cdot[Y_{i,t}-Y_{i,t-1}-\hat{m}_{d^{\prime}_{t}}(t,t-1,\mathbf{D}_{i,t-1},X_{i})]/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D}_{i,t-1},X_{i})}{(\sum_{j=1}^{n}\omega(D_{j,t};d^{\prime}_{t},h))\cdot\hat{\rho}_{d_{t}}(\mathbf{D}_{i,t-1},X_{i})/\hat{\rho}_{d^{\prime}_{t}}(\mathbf{D}_{i,t-1},X_{i})}\right].$
Furthermore, the algorithm can be easily modified to estimate $\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}]$, the ATET of a lagged treatment dose on a later outcome as defined in (10), by using of the sample analogs of DR expressions (3) or (20) in the case of repeated cross-sections or panel data, respectively.
For the variances of our estimators, we adapt the asymptotic variance approximation suggested in Zhang (2025) to our settings with non-zero treatment doses in both the treatment and control groups. Considering the case of repeated cross-sections, we define
@@
with $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E\left[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)\right]$ being the smoothed ATET with bandwidth $h$. Then $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta):=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the smoothed Neyman-orthogonal score function for $\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ with bandwidth $h$. The asymptotic variance, denoted by $\sigma_{h}^{2}$, of our DML estimator of $\Delta_{d_{t},d_{t}^{\prime},t}$ corresponds to
It is worth noting that the second term in the asymptotic variance in equation (25)
@@
arises from a Taylor expansion of the score w.r.t. the (low dimensional) density $P(D_{t}=d_{t},T=t)$.
A cross-fitted variance estimator can be constructed based on the sample analog of equation (25):
$\hat{\sigma}_{h}^{2}=\frac{1}{K}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left[\left(\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta}_{k})-\frac{\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}}{\hat{\Pi}_{k,d_{t},t}}\left(\omega(D;d,h)\cdot I\{T=t\}-\hat{\Pi}_{k,d_{t},t}\right)\right)^{2}\right],$ (26)
where $\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h}$ is the cross-fitted ATET estimate (see step 4 of the algorithm). The term $\phi^{h}(W_{i},\hat{\Delta}_{d_{t},d_{t}^{\prime},t}^{h},\hat{\Pi}_{k,d_{t},t},\hat{\eta}_{k})$ denotes the estimate of score function $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},\Pi_{d_{t},t},\eta)$ for subject $i$, which is computed using the subsample estimates of the nuisance parameters $\hat{\eta}_{k}$ (see step 3 of the algorithm). Similarly $\hat{\Pi}_{k,d_{t},t}=|\mathcal{W}_{k}^{C}|^{-1}\sum_{i}\omega(D_{i};d,h)\cdot I\{T_{i}=t\}$ is the subsample estimate of the density $\Pi_{d_{t},t}$. In Zhang (2025), such types of variance estimators are shown to be consistent, provided that the kernel bandwidth $h$ satisfies $h^{-2}\varepsilon_{n}^{2}+h^{-3}n^{-1}=o(1)$ for $\varepsilon_{n}=o(n^{-1/4})$.
@@
where $\Delta_{d_{t},d_{t}^{\prime},t}^{h}=E[\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)]$ is the smoothed ATET under bandwidth $h$ and $\phi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)=\psi^{h}(W,\Delta_{d_{t},d_{t}^{\prime},t}^{h},P_{d_{t}},\eta)-\Delta_{d_{t},d_{t}^{\prime},t}^{h}$ is the score function, with